# Why These Documents? Explainable Generative Retrieval with Hierarchical Category Paths

### v1

@@ 5.2. Explainability of HyPE (RQ1 & RQ2)
As shown in Table 1, applying HyPE improves overall semantic relevance across all baselines. This indicates that HyPE’s category path effectively captures and explains the relationship between the query and the document. Notably, HyPE achieves higher overall relevance than the term-matching method (i.e., BM25), further proving the validity of the category path as an explanation.
Table 3. Retrieval accuracy of existing generative retrieval methods and HyPE framework on the MS MARCO dataset. * denotes the statistical significance on paired t-test $p<0.05$.
Method MS MARCO
R@1 R@10 R@100 M@10
NOVO (Wang et al., 2023) 49.1 80.8 92.5 59.2
[…]
+ HyPE 32.2 62.7 78.5 41.9
Improvement +1.6% +2.5% +1.7% +2.2%
SEAL (Bevilacqua et al., 2022) 25.9 68.6 87.9 40.2
[…]
+ HyPE 28.4 57.5 73.1 37.8
Improvement +1.1% +3.6% +2.2% +2.7%
GENRET (Sun et al., 2023) 47.9 80.8 91.6 59.2
[…]
+ HyPE 44.9 74.6 87.1 54.7
Improvement +2.3% +1.4% +1.8% +1.7%
5.2.2. Qualitative analysis (RQ2).
We evaluate the quality of explanations generated by HyPE through a human evaluation conducted via Amazon Mechanical Turk (AMT). We randomly sample 100 examples from the NQ320K dev set and ask three human judges to compare the quality of the explanations based on four distinct criteria: overall, specificity, reasonability and comprehensiveness. Detailed descriptions of these criteria are presented in Appendix A.3.
In Figure 4, HyPE outperforms both the title docid baseline and BM25 across all criteria, receiving high scores for its overall explanation of the retrieval process. Specifically, HyPE shows substantial margin of superiority in terms of specificity and reasonability. This demonstrates that HyPE provides clearer explanations of retrieval process, as well as more logical and reasonable explanation. Furthermore, HyPE beats other baselines in comprehensiveness, indicating that its hierarchical category path is effective in explaining not only narrow, specific details but also broader semantic information. These results highlight that HyPE’s pseudo-reasoning, which utilizes hierarchical category paths, provides users with a effective and reasonable explanation of the retrieval process.
Following human evaluation, we conduct two types of case studies to verify our HyPE provides more effective explanations for each query. Figure 3 illustrates HyPE’s pseudo-reasoning process in cases where a single document is annotated with multiple queries on different topics. In Example 1, for the query “the core of the sun in which the sun’s thermonuclear energy is produced”, the model generates paths related to the universe and energy conversion, clearly explaining the thematic relevance between the query and the document. However, for another query, “what stage of the star life cycle is the sun in”, it generates a path related to stellar evolution, which is different from the previously generated path but relevant to the query. This shows that HyPE can provide effective explanations to users by tailoring each query. Moreover, in Example 2, HyPE generates the same path for different queries with similar topics, specifically related to the building of the Statue of Liberty. This highlights HyPE ’s capability to provide coherent and consistent explanations for queries on related topics.
Figure 5 illustrates the difference between using only the title as the docid and considering paths alongside it with HyPE. For the query “Who sings love me tender in princess diaries 2?”, the baseline model using only the title fail to fully understand the query and retrieve an incorrect document [The Princess Diaries 2: Royal Engagement]. In contrast, when HyPE is adopted, the generative retrieval model first generates the path “entertainment ¿ theatre ¿ theatre soundtracks”, which captures the context related to “sing” that is missed by relying solely on the title. This allows the model to understand the query more specifically and successfully retrieve the correct document [The Princess Diaries soundtrack].
5.3. Retrieval Accuracy of HyPE (RQ3)
Table 2 shows retrieval accuracy of various docid types with HyPE on NQ320K, comparing the results against each baseline. Overall, HyPE consistently improves retrieval accuracy across all docid types in both seen test and unseen test. This demonstrates that HyPE’s hierarchical category paths can be orthogonally applied to enhance retrieval accuracy across different docid types, suggesting that integrating these paths into existing generative retrieval methods can further improve performance. While HyPE can be applied to all docid types effectively, the experimental results show that title docid yields the most significant performance improvement when HyPE is applied. Our paths, serve as a pseudo-reasoning allowing the model to navigate step-by-step through various semantic hierarchical categories before arriving docid. Since title docids are concise and inherently reflect a structured overview of a document, they aligns well with the HyPE’s hierarchical category paths, further enhancing retrieval accuracy and explanation clarity.
Figure 6. Performance changes of HyPE w.r.t. the number of decoded category paths to obtain a ranked docid list.
Additionally, we also report the experimental results on MS MARCO in Table 3. In contrast to NQ320K, MS MARCO, as a web search-based dataset, has varying document formats and contains significant semantic noise. Despite these challenges, the results show that HyPE consistently improves retrieval accuracy on this dataset as well. Although the hierarchical category paths are constructed using Wikipedia categories as the backbone, the consistent performance gains on MS MARCO emphasize the robustness and generalizability of HyPE. These findings suggest that HyPE can be widely applied to datasets across various domains in the future.
5.4. Analysis of Path-Aware Reranking (RQ4)
To validate the effectiveness of path-aware reranking strategy, we analyze the performance changes in retrieval accuracy with respect to the number of hierarchical category paths considered by HyPE. Figure 6 presents the analysis results, showing that retrieval accuracy improves as the number of paths increases across all docid types. Notably, there is a clear performance gap between the setting without path-aware reranking strategy (i.e., $K=1$) and with path-aware reranking strategy (i.e., $K>2$). These results strongly indicate that considering multiple paths through the path-aware reranking strategy allows the most relevant docids to be prioritized in the final ranked list, thereby enhancing retrieval accuracy. However, we observe that using too many paths eventually leads to a plateau in performance improvement. Beyond a certain threshold, additional paths tend to introduce noise or increase unnecessary computational complexity. Through our experiments, we can conclude that using three paths yields the most optimal retrieval accuracy for the majority of docid types.
6. Conclusion
This paper proposes HyPE, a novel framework designed to enhance the explainability of document retrieval by utilizing hierarchical category paths. By leveraging hierarchical category paths, HyPE addresses the limitations of existing generative retrieval methods that struggle with providing meaningful explanations for why specific documents are retrieved. Furthermore, we introduce path-aware reranking allows for the aggregation of diverse topic information, ensuring that the most relevant docids are prioritized in the final ranked list of docids. Our extensive experiments demonstrate that HyPE not only provides reliable and reasonable explanations but also enhances overall retrieval performance. To the best of our knowledge, this study is the first attempt to enhance explainability in generative retrieval, and we hope our research paves the way for positive advancements in this field.
Appendix A Appendix
A.1. Backbone category hierarchy
To address the criteria mentioned in Section 4.1—Semantic Hierarchy, Generalizability, and Specificity—we utilize Wikipedia’s category tree as the foundation for our hierarchical structure, designating the Main Topic classification category as the root node of the hierarchy. The Main Topic classification category is chosen as the root because it offers a broad and inclusive starting point, covering diverse domains. From this root node, our hierarchy branches into 40 child categories, each representing different fields such as Economy, Language, and Science, among others. These child categories, in turn, contain more specific subcategories that further detail their respective domains. To utilize Wikipedia’s category tree, we employed Selenium [^10] https://pypi.org/project/selenium/ to recursively scrape the Wikipedia and extract the Wikipedia category tree. When linearizing the category hierarchy into a hierarchical category path, each category is connected using the delimiter ¿. The delimiter ¿ is chosen among several candidate delimiters because it showed the highest semantic similarity to the natural language sentence “the right category is included in the left category”, as measured by Sentence-T5.
@@ A.1. Backbone category hierarchy
Table 4. Statistics of the used category hierarchy, showing the number of nodes at each level (or depth).
A.2. Dataset overview
In this work, we use NQ320K and MS MARCO. For NQ320K, we follow NCI (Wang et al., 2022) setup and adhered to the seen and unseen test splits used in GENRET (Sun et al., 2023). For MS MARCO, we construct dataset based on the MSMARCO document ranking dataset, following setups from Ultron (Zhou et al., 2022), GENRET (Sun et al., 2023), and NOVO (Wang et al., 2023). Methods using the MSMARCO passage ranking dataset, such as TOME (Ren et al., 2023) and GLEN (Lee et al., 2023), or using NQ100K like SE-DSI (Tang et al., 2023), are excluded from consideration as reference baselines unless they report results specifically on the our dataset setup.
@@ A.2. Dataset overview
Table 5. Statistics of the document retrieval datasets used.
A.3. Human Evaluation
We assess the quality of the generated explanations by conducting a human evaluation, where we compare the outputs of HyPE to other baseline models using Amazon Mechanical Turk (AMT). We show the interface for the evaluation in Figure 7. We ask three human judges to evaluate the explanations based on the following four criteria.
  • Overall Which retrieval system output better explains the retrieval process overall?
  • Specificity Which retrieval system output provides more specific information?
  • Reasonability Which retrieval system output represents the retrieval process more logically and reasonably?
  • Comprehensiveness Which retrieval system output more comprehensively reflects the content of the document?
A.4. Prompts
@@
Figure 7. Annotator interface of human evaluation on retrieval system output.
Table 6. The prompt for building final candidate path set.
@@
You’re a taxonomy expert.
You will receive a document along with a set of candidate taxonomy hierarchy paths for the document.
Your task is to select the path that can represent the document. Exclude paths that are too broad or less relevant or contain too specific information such as year.
@@
¡Document title¿
{Document title}
¡Document contents¿
{Document contents}
¡Candidate taxonomy hierarchy paths¿
{pre-candidate path set}
¡Selected representative taxonomy hierarchy paths¿
{Candidate path set}

### v2

@@ A.1 Quantitative Analysis of Explainability
As shown in Table 6, applying HyPE improves overall semantic relevance across all baselines. This indicates that HyPE’s category path effectively captures and explains the relationship between the query and the document. We note that HyPE achieves higher overall relevance than the term-matching method (i.e., BM25), further proving the validity of the HyPE’s category path as an explanation. Moreover, maximum level of hierarchical category path significantly influences overall semantic relevance. Specifically, paths with fewer levels than the default level (level 4) fail to capture sufficient semantic relevance between the query and the document, resulting in limited explainability. These results demonstrate that for category paths to effectively serve as explanations, they must achieve specificity necessary to sufficiently explain specific and detailed semantic information, as mentioned in Section 3.1.
A.2 Pseudo-Reasoning
Generating the hierarchical path resembles step-by-step reasoning. However, unlike natural language-based reasoning in LLM, we use the term “pseudo-reasoning” because the path structure is more akin to pseudo-code.
A.3 Backbone category hierarchy
Criteria for Selecting the Backbone.
To address the criteria mentioned in Section 3.1—Semantic Hierarchy, Generalizability, and Specificity—we utilize Wikipedia’s category tree as the foundation for our hierarchical structure, designating the Main Topic classification category as the root node of the hierarchy.
  • Specificity: Are they capable of sufficiently explaining specific and detailed information?
@@ A.3 Backbone category hierarchy
Table 7: Statistics of the used category hierarchy, showing the number of nodes at each level (or depth).
Wikipedia category tree Overview.
Wikipedia’s category tree consists of 40 nodes at level 1, covering broad categories such as Business, Sports, Science, Philosophy, Language, Health, Government, Culture, and others. This feature of encompassing a wide range of fields ensures that Wikipedia’s category tree satisfies the criterion of Generalizability, as it can be applied across various domains. Moreover, these broad categories are further subdivided into increasingly specific subcategories as the level increases. For instance, level 1 Science is divided into major subcategories such as Branches of Science, Scientists, and History of Science at level 2. Among these, Branches of Science is further refined into Applied Science, Formal Science, and Social Science at level 3, which are then expanded into even more specific subcategories like Computer Science, Agronomy, Metrology, and Bioinformatics at level 4. As the levels progress, the structure captures increasingly detailed semantic information, effectively fulfilling the criterion of Specificity. Additionally, the broad-to-specific hierarchical structure of Wikipedia’s category tree naturally achieves Semantic Hierarchy.
Implementation Details for Path.
To utilize Wikipedia’s category tree, we employed Selenium [^4] https://pypi.org/project/selenium/ to recursively scrape the Wikipedia and extract the Wikipedia category tree. When linearizing the category hierarchy into a hierarchical category path, each category is connected using the delimiter >. The delimiter > is chosen among several candidate delimiters because it showed the highest semantic similarity to the natural language sentence “the right category is included in the left category”, as measured by Sentence-T5.
Scalability of Our Backbone Hierarchy.
We believe that Wikipedia’s category tree will function effectively in most document retrieval scenarios. This taxonomy was specifically designed to systematically categorize real Wikipedia documents, which cover a wide range of domains and knowledge. Its broad and deep structure ensures that it can encompass diverse domains effectively, making it a strong backbone hierarchy for general-purpose retrieval systems.
Adaptability of HyPE.
However, we acknowledge that in more specialized domains—such as expert-driven fields like medicine, law, or scientific literature—the Wikipedia-based hierarchy may not fully capture domain-specific semantics or categorization needs. In such cases, the backbone hierarchy may need to be replaced or augmented with a domain-specific taxonomy better suited to the task. We note that HyPE is compatible with this setting: domain-specific taxonomies can be integrated in a plug-and-play fashion. For example, the domain taxonomy used for academic paper retrieval Kang et al. (2024b) could be adopted as an alternative backbone in that context. Furthermore, if a well-defined taxonomy does not yet exist for a specific domain, one can be constructed using taxonomy induction methods Zhang et al. (2018); Lee et al. (2022).
A.4 Dataset Overview
@@
Table 8: Statistics of the document retrieval datasets used.
In this work, we use NQ320K and MS MARCO. For NQ320K, we follow NCI Wang et al. (2022) setup and adhered to the seen and unseen test splits used in GENRET Sun et al. (2023). For MS MARCO, we construct dataset based on the MSMARCO document ranking dataset, following setups from Ultron Zhou et al. (2022), GENRET Sun et al. (2023), and NOVO Wang et al. (2023). Table 8 shows the statistical details of the datasets used in our experiments.
A.5 Human Evaluation
We assess the quality of the generated explanations by conducting a human evaluation, where we compare the outputs of HyPE to other baseline models using Amazon Mechanical Turk (AMT). In this experiment, we use the title docid baseline described in Section 4.1, and additionally include BM25 as a baseline. which is capable of providing explanations for its retrieval results by highlighting the top-ranked terms contributing to the retrieval. We ask human judges to evaluate each sample’s explanations based on the following four criteria.
  • Overall: Which retrieval system output better explains the retrieval process overall?
  • Specificity: Which retrieval system output provides more specific information?
  • Reasonability: Which retrieval system output represents the retrieval process more logically and reasonably?
  • Comprehensiveness: Which retrieval system output more comprehensively reflects the content of the document?
Note that our human evaluation involved a total of 300 human judges, with each sample being independently evaluated by 3 different human judges. This setting is designed by referencing previous works that conduct human evaluation Kim et al. (2023); Lee et al. (2025). We show the interface for the human evaluation in Figure 5
A.6 Human Reranking
To evaluate whether explanations provided by HyPE can help users more effectively identify relevant documents in realistic search scenarios, we conduct a human reranking experiment via Amazon Mechanical Turk (AMT). We prepare two conditions for comparison: (1) a title-only setting and (2) a title+path setting, where the title is shown along with a hierarchical category path explanation generated by HyPE. For each query, five candidate documents are shown in both conditions, with the same title across settings; only the presence or absence of the category path differs, allowing for a controlled comparison of explanation impact. We randomly sample 100 query-document instances from the NQ320K dev set where the title docid baseline with HyPE successfully retrieves the gold document within the top-5 results. Human judges are asked to (1) rank the five candidates based on their relevance to the query (i.e., human reranking), and (2) indicate their confidence in the ranked list using a 5-point Likert scale. Based on the collected responses, we compute three metrics: Recall@1, which indicates whether the gold document was ranked first; MRR@5, which reflects how highly the gold document was ranked; and Confidence, which measures how certain participants are in their rankings. This setup allows us to quantitatively assess whether the explanations produced by HyPE improve both the accuracy and certainty of user decisions in realistic, information-limited search environments. We show the interface for the human reranking in Figure 6
A.7 Implementation Details
We use T5-base Raffel et al. (2020) as our backbone model. For the input of the indexing task, we utilize the FirstP approach as our document representations (Section 2). Additionally, for the indexing task, we employ five synthetic queries, generated by using docT5query Nogueira and Lin (2020) with nucleus sampling with parameters $p=0.8$ and $t=0.8$. We use new [DOC] token to separate the path from the docid, which we insert between the path and the docid. We optimize our model as described in 3.2, while employing AdamW optimizer with a learning rate of 5e-4 and a batch size of 128, for up to 1M training steps. During the inference of HyPE, we adopt path-aware ranking strategy; for the path generation stage, we generate three category paths (i.e., $K_{p}=3$), and for the docid generation stage, we use constrained beam search with a beam size of 100 (i.e., $m=100$). To build the summary docid baseline and keyword docid baseline, we utilize the off-the-shelf text summarization model based on BART Lewis et al. (2020) and the keyword extraction tool Grootendorst (2020).
A.8 Analysis of Efficiency
To quantify the inference cost introduced by generating hierarchical category paths, we measure the average inference time per instance using an NVIDIA RTX 4090 GPU. Specifically, we compare two decoding settings: (1) decoding only the docid, and (2) decoding both the docid and a single hierarchical category path. Our results show that the additional decoding required for generating a single path introduces only a marginal increase in inference time, demonstrating that HyPE ’s explainability can be achieved with minimal efficiency loss.
A.9 Prompt
@@
You’re a taxonomy expert. You will receive a document along with a set of candidate taxonomy hierarchy paths for the document. Your task is to select the path that can represent the document. Exclude paths that are too broad or less relevant or contain too specific information such as year.
@@
<Document title>: {Document title}
<Document contents>: {Document contents}
<Candidate hierarchy paths>: {pre-candidate path set}
<Selected hierarchy paths>: {Candidate path set}
Table 9: The prompt for building final candidate path set.
Figure 5: Annotator interface of human evaluation on retrieval system output.
Figure 6: Annotator interface of human reranking on retrieval system output.